# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

### v1

@@ Abstract.
Abstract
Client-Service Representatives (CSRs) are vital to organizations. Frequent interactions with disgruntled clients, however, disrupt their mental well-being. To help CSRs regulate their emotions while interacting with uncivil clients, we designed Pro-Pilot, an LLM-powered assistant, and evaluated its efficacy, perception, and use. Our comparative analyses between 665 human and Pro-Pilot-generated support messages demonstrate Pro-Pilot’s ability to adapt to and demonstrate empathy in various incivility incidents. Additionally, 143 CSRs assessed Pro-Pilot’s empathy as more sincere and actionable than human messages. Finally, we interviewed 20 CSRs who interacted with Pro-Pilot in a simulation exercise. They reported that Pro-Pilot helped them avoid negative thinking, recenter thoughts, and humanize clients; showing potential for bridging gaps in coworker support. Yet, they also noted deployment challenges and emphasized the irreplaceability of shared experiences. We discuss future designs and societal implications of AI-mediated emotional labor, underscoring empathy as a critical function for AI assistants in front-office roles.
@@ 1. Introduction
When we engage an organization for their product or service, our initial contact is with staff known as front-office workers or Client Service Representatives (CSRs). They are the first line of response for the organization. Unlike other roles within an organization, a CSR’s task involves frequent interactions with clients and individuals outside an organization (Hochschild, 1983). These interactions require CSRs to manage their emotions constantly to complete tasks. Essentially they exert emotional labor to appear professional (Hochschild, 1983). The crucial challenge for a CSR is engaging with a client who starts behaving uncivilly by communicating in a rude, aggressive, and emotionally charged way (Grandey et al., 2007). No matter the type of request, a CSR’s role is to resolve a client’s concern and comply with the adage, “the customer is always right.” Unfortunately, this leads to an emotional dissonance between what a CSR expresses and what they actually feel (Grandey, 2000). Ultimately, such workers are vulnerable to excessive stress and eventual burnout. Not only do CSRs report being emotionally depleted and detached, but they also report a lack of accomplishment (Brotheridge and Grandey, 2002). Clearly, CSRs play a critical role within the organization, but, we have witnessed little innovation in alleviating their emotional toll. Our paper investigates how AI-coworkers help CSRs regulate their emotions in the face of client incivility.
The brunt of client incivility in front-office work makes it infamous for low satisfaction and high-turnover (Pienaar and Willemse, 2008). A fundamental solution to the emotional distress of this role is Emotional Regulation (ER) (Yang and Lau, 2019). Basically, ER is the process through which one rethinks a negative situation (Grandey, 2000). While a worker may be able to do this on their own, research shows that coworkers play an important role in supporting ER (Yang and Lau, 2019). A good coworker can read the emotional cues of CSR’s work tasks and provide suggestions to help minimize the brunt of an aggressive client. However, CSRs are increasingly adopting remote work setups (Haan and Holznienkemper, 2023), which dampens social support (Vacchiano et al., 2024). Meanwhile, organizational scientists are calling for digital interventions to support worker wellbeing at scale (Black et al., 2019). We answer this call by designing and evaluating, Pro-Pilot— a Large Language Model (LLM)–based AI assistant for on-task Emotional Regulation in front-office work. While generative AI is emerging as a potent tool to complement the informational load of many different roles, the HCI community lacks research to investigate their use in emotional labor.
@@
We developed Pro-Pilot by leveraging domain knowledge on client incivility (Andersson and Pearson, 1999; Frey-Cordes et al., 2020; Cho et al., 2002), real-world complaint data (Axelbrooke, 2017), and recent advancements in LLM-powered cognitive change (Burger et al., 2021; Sharma et al., 2023) and simulation (Shaikh et al., 2023). To answer the first research question, we conducted data-generation and evaluation tasks with 259 CSRs to comparatively analyze Pro-Pilot and human-coworker support for a variety of client incivility situations (1(a)). To answer the second question, we used Pro-Pilot as a technology probe and conducted a mixed-methods simulation exercise to understand how real CSRs could include it in their client interactions by juxtaposing Pro-Pilot usage with CSR’s socio-organizational norms (1(b)). Consequently, we contribute:
@@
  • Pro-Pilot: an interactive technological artifact to expose CSRs to client incivility and learn healthy, long-term emotional labor practices to improve their own health and support their coworkers.
@@
  • Empirical evidence that LLM–based empathetic support can be engineered to adapt to— and express empathy in— various client-incivility scenarios (RQI). Our results demonstrate that Pro-Pilot’s messages are linguistically distinct from both zero-shot approaches and human–coworkers, and moreover, Pro-Pilot’s messages were perceived to be more empathetic on several dimensions including sincerity, actionability, and relatability.
@@
  • End-user insight on the function of LLM–based empathetic support to scaffold them through emotional labor during uncivil interactions (RQ2). Our findings showcase Pro-Pilot’s process of redirecting negativity and making CSRs more self-efficacious. While CSRs envision Pro-Pilot addresses important opportunities in workplace social support, they also surfaced the challenges of Pro-Pilot emulating holistic human support.
@@
This paper has implications for re-imagining how AI-assistants for workers should be designed and also re-imagining the social norms and policies to accommodate these advancements.
@@
Reflexive Considerations. Front-office work has many stakeholders, including the employer and the clients. However, it is the CSR who bears the burden of repeated emotional labor (Grandey, 2000). The relationship between these stakeholders is asymmetric, as employers can replace personnel and clients can switch services, but the CSR does not possess the same mobility (Fisk and Neville, 2011). Following from recent works that take a worker-centered perspective (Das Swain et al., 2023a, 2024; Kawakami et al., 2023), our research aims to illuminate the worker who is disadvantaged. This paper focuses on the needs of CSRs and centers their perspective throughout the evaluation. Two authors have past experience in front-office roles and direct end-user servicing. Four authors are researchers in an organization that employs its own CSRs. They helped us access real CSRs (and their resources) to provide feedback on our study design. We recruited participants for this study outside their organization to capture perspectives from different organizational sectors. All evaluations described in the paper were approved by the IRB of the first author’s institute.
@@
(a) Technical evaluation: Compared Pro-Pilot’s support messages with those produced by human-coworkers in CSR roles.
@@
(b) User evaluation: Studied participant experiences with Pro-Pilot’s emotional support while interacting with uncivil clients.
@@ 2.1. Front-Office Work and Incivility
People who work front-office are the human touch-points for direct interaction with clients  [^1] The literature uses the terms “client” and “customer” interchangeably. We use the term “client” to refer to anyone who purchases or avails a good or service.. The end-user, or client, can directly observe and experience the activities of front-office workers (Zomerdijk and de Vries, 2007). Front-office work spans a variety of organizational sectors including health (e.g., front-desk), hospitality (e.g., air staff), retail (e.g., store associates), and technology (e.g., customer service). For front-office roles, or “pink-collar” work, organizations seek to hire individuals who can emotionally invest in their clients, meet their needs and please them (Hochschild, 2019). Our research focuses on online front-office work where the employee frequently interacts with their client via text-based communication. This paper refers to these workers as Client Support Representatives or (CSRs).
Today, CSRs cater to a variety of clients and it is common to encounter high-maintenance clients who have unreasonable service demands (Fisk and Neville, 2011). Unfortunately, these demands translate into rude and discourteous behaviors. Koopmann et al. defines client incivility as “low-quality interpersonal treatment that employees receive during service interactions.” Unlike overt verbal abuse, such as calling the CSR names and expletives, or downright violence, incivility is more implicit. According to Andersson and Pearson, the intent to harm is ambiguous and can be deflected, e.g., “I didn’t mean to be rude; I was just in a hurry” (Andersson and Pearson, 1999). Organizational psychologists have noted a possible reason for incivility to be the increase in client entitlement while seeking services (Yagil, 2008) and a lack of consideration for others (Pearson et al., 2000). Meanwhile, a CSR role is often low-wage, considered low-skill, and lacks the decision-latitude needed to respond to mistreatment from clients (Fisk and Neville, 2011). Coupled together, incivility creates a dynamic between the client and CSR, where the latter feels injustice and negative emotion (Fisk and Neville, 2011). Therefore, a CSR’s job involves emotional labor because of two processes; managing the interpersonal demands of interaction and controlling their own negative emotions (Grandey, 2000). Facing client incivility regularly puts a CSR at the risk of experiencing emotional exhaustion (lack of motivation and focus), depersonalization (detachment from others’ emotions), and reduced accomplishment (feeling of low effectiveness) (Maslach et al., 1997). The reduced emotional health of CSRs affects organizations as well. They risk an increased turnover and rise in negative work attitudes (Brotheridge and Grandey, 2002). Given the large volume of workers in this role, we need to urgently design solutions for their wellbeing.
@@ 2.2. Social Support and Emotion Regulation
At work, social support is known to be an important moderator for incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012; Xiao et al., 2022). One kind of social support is emotional support, where a coworker provides sympathy and understanding (Caplan, 1975). This form of support can help reduce work-related stress (House, 1983). Studies have shown that CSRs with more supportive coworkers are able to recover from the negative effects of client incivility, by being able to recognize emotions and guide constructive actions in response (Xu et al., 2014; Wong and Law, 2017). Yet, in practice, CSRs are known to receive low levels of emotional support (Walsh and Deery, 2006) as coworkers can favor client interests over that of an employee (Yang and Lau, 2019). A coworker with high emotional intelligence is someone capable of expressing compassion in a way that helps the CSR cope with negative emotions and proceed to their goal (Law et al., 2004). This behavior of coworkers was the primary metaphor we used to conceive Pro-Pilot.
@@
Emotional support can be viewed as a form of emotional coping, or Emotional Regulation (ER). Unlike traditional discussions of ER, coworker support is other-directed (Yang and Lau, 2019). Conversely, CSRs are trained to and expected to regulate their emotions using Surface Acting — they adjust the emotion they present to clients (Grandey, 2000). However, even though Surface Acting is a requirement for roles with emotional labor, suppressing emotions at work can lead to a reduction in job satisfaction (Côté and Morgan, 2002). Instead, Grandey recommends CSRs to engage in Deep Acting — they adjust their perception of the situation. Deep Acting has many strategies, but, CSR’s have limited job mobility and do not have the decision latitude to change their clients (Grandey, 2000). As a result, we designed Pro-Pilot’s core component Emo-Reframe to achieve ER by helping the CSR cognitively reappraise the negative interaction.
@@
HCI research has explored ER along several different dimensions. However, Slovak et al. found that most studies have focused on suppression of emotions (Slovak et al., 2023), and in the CSR context, these methods will have the same limitations as Surface Acting (Côté and Morgan, 2002). Hence, we pursue the ER strategy of cognitive reappraisal or cognitive change. Although, cognitive reappraisal for ER has been explored in HCI in the form of therapy (such as CBT) (Sharma et al., 2023), we find scant evidence for delivering ER in-context and guiding users through the process (Slovak et al., 2023). These gaps motivated us to think beyond training modules and post-hoc support, to implement Pro-Pilot for on-task ER within the CSR’s workflow.
2.3. HCI and AI for Emotion Labor
@@
A worker performs emotional labor when their job role expects them to either maintain certain emotions or evoke certain emotions (Hochschild, 1983). The HCI community has primarily studied this phenomenon in the scope of crowd work (Menking and Erickson, 2015) and particularly content moderation (Wohn, 2019; Dosono and Semaan, 2019). Online communities rely on volunteer moderation, and users who participate as moderators often work in a way analogous to front-office work (Section 2.1). Even in an online — primarily text-based — interaction, content moderators have described adopting emotion management techniques to tackle their tasks (e.g., receiving threatening messages from users) (Dosono and Semaan, 2019). Content moderators differ from CSRs, as they are less likely to have synchronous interactions. Having said that, these workers are also subject to reduced emotional wellbeing, such as lack of appreciation and negativity (Wohn, 2019). Another key distinction between content moderation and CSRs is the context of employment. The former is often a voluntary role, whereas the latter is likely to be one’s primary employment. These differences significantly change the socio-organizational dynamics between the worker–client and among the workers. The unique normative structure of emotional labor for CSRs motivated us to conduct a user evaluation to inspect Pro-Pilot with actual users.
@@
Since the emotional labor in content moderation is largely composed of harmful content and secondary trauma, solutions have focused on dynamic content filtering and modulation (Das et al., 2020; Karunakaran and Ramakrishan, 2019). These approaches of preventing or reducing exposure (Steiger et al., 2021) are incompatible with CSR roles where the emotional labor is not caused by explicitly harmful content but by incivility, which is inherently ambiguous, and originates from the client (Andersson and Pearson, 1999).  Cook et al.’s approach to inject positive stimuli in between tasks could be compatible, but their effectiveness was mixed (Cook et al., 2022). Alternatively, Osawa suggested an on-task solution for face-to-face emotion labor where enhanced glasses make the user’s eyes appear to be emotionally invested (Osawa, 2014). The equivalent of this type of substitution of emotional labor would be the introduction of AI-powered chatbots. Arguably, these chatbots have reinforced client- interactions (Sweezey, 2019), but on one hand, chatbots relieve the task burden of routine inquiries; on the other hand, a large proportion of clients still prefer human CSRs for more complex complaints and better emotional support (Elliott, 2018). Thus conversational agents do not sufficiently mitigate the challenges of emotional labor on CSR.
@@
3. Pro-Pilot: System Description
In traditional front-office work, coworkers are an important social resource. When a CSR encounters uncivil clients, they often consult their coworkers to assuage the excessive emotional labor (Xu et al., 2014; Fisher, 2014). Such consultations can be challenging to interleave into the ongoing task of addressing an aggressive client. Thus, we used OpenAI’s GPT-4o (gpt, 2024) to design an emotional support utility that can be embedded into the CSR’s task environment. Our utility differs from typical conceptions of AI-assistants for work, or “copilots,” that provide informational support by enabling better problem solving (Cutrona and Russell, 1987), such as coding assistance (Mozannar et al., 2024). Instead, our utility leverages LLMs to provide emotional support by enabling ER through cognitive change (Grandey, 2000). Throughout the paper, we refer to our overarching system for empathetic support as Pro-Pilot. This section explains how Pro-Pilot was implemented. First, we describe how we harnessed domain knowledge and real complaint datasets to generate uncivil client interactions. Second, we describe how Pro-Pilot produces empathetic messages for CSRs who are involved in uncivil interactions. Later, in Section 4.4, we compare how Pro-Pilot performs against other state-of-the-art LLMs (GPT, Llama, Mistral), which additionally justifies our approach for designing and fine-tuning Pro-Pilot for the purposes of our study.

### v2

@@ Abstract.
Client-Service Representatives (CSRs) are vital to organizations. Frequent interactions with disgruntled clients, however, disrupt their mental well-being. To help CSRs regulate their emotions while interacting with uncivil clients, we designed Care-Pilot, an LLM-powered assistant, and evaluated its efficacy, perception, and use. Our comparative analyses between 665 human and Care-Pilot-generated support messages highlight Care-Pilot’s ability to adapt to and demonstrate empathy in various incivility incidents. Additionally, 143 CSRs assessed Care-Pilot’s empathy as more sincere and actionable than human messages. Finally, we interviewed 20 CSRs who interacted with Care-Pilot in a simulation exercise. They reported that Care-Pilot helped them avoid negative thinking, recenter thoughts, and humanize clients; showing potential for bridging gaps in coworker support. Yet, they also noted deployment challenges and emphasized the indispensability of shared experiences. We discuss future designs and societal implications of AI-mediated emotional labor, underscoring empathy as a critical function for AI assistants for worker mental health.
large language models, empathy, emotion regulation, emotional labor, human-AI interaction, future of work, mental health
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713705isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Collaborative and social computing systems and toolsccs: Applied computing Psychologyccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
When we engage with an organization for their product or service, our initial contact is typically with staff known as front-office workers or Client Service Representatives (CSRs). They are the first line of response for the organization. Unlike other roles within an organization, a CSR’s task involves frequent interactions with clients and individuals outside an organization (Hochschild, 1983). These interactions require CSRs to constantly manage their emotions to complete tasks. Essentially, they exert emotional labor to appear professional (Hochschild, 1983). The crucial challenge for a CSR arises when engaging with a client who starts behaving uncivilly by communicating in a rude, aggressive, and emotionally charged manner (Grandey et al., 2007). No matter the type of request, a CSR’s role is to resolve a client’s concern and comply with the adage, “the customer is always right.” Unfortunately, this leads to an emotional dissonance between what a CSR expresses and what they actually feel (Grandey, 2000). Ultimately, such workers are vulnerable to excessive stress and eventual burnout. Not only do CSRs report being emotionally depleted and detached, but they also report a lack of accomplishment (Brotheridge and Grandey, 2002). Clearly, CSRs play a critical role within the organization, but, we have witnessed little innovation in alleviating their emotional toll. Our paper investigates how AI-coworkers help CSRs regulate their emotions in the face of client incivility.
The brunt of client incivility in front-office work makes it notorious for low satisfaction and high-turnover (Pienaar and Willemse, 2008). A fundamental solution to the emotional distress of this role is Emotional Regulation (ER) (Yang and Lau, 2019). Basically, ER is the process through which one rethinks a negative situation (Grandey, 2000). While a worker may be able to do this on their own, research shows that coworkers play an important role in supporting ER (Yang and Lau, 2019). A good coworker can read the emotional cues of CSR’s work tasks and provide suggestions to help minimize the brunt of an aggressive client. However, CSRs are increasingly adopting remote work setups (Haan and Holznienkemper, 2023), which dampens social support (Vacchiano et al., 2024). Meanwhile, organizational scientists are calling for digital interventions to support worker wellbeing at scale (Black et al., 2019). We answer this call by designing and evaluating Care-Pilot— a Large Language Model (LLM)–based AI assistant for on-task Emotional Regulation in front-office work. While generative AI is emerging as a potent tool to complement the informational load of different roles, the HCI community lacks research to investigate their use in emotional labor.
@@
The paper is organized into three key sections. First, Section 3: System Description details how we developed Care-Pilot. We cover our approach to leverage domain knowledge on client incivility (Andersson and Pearson, 1999; Frey-Cordes et al., 2020; Cho et al., 2002), real-world complaint data (Axelbrooke, 2017), and recent advancements in LLM-powered cognitive change (Burger et al., 2021; Sharma et al., 2023) and simulation (Shaikh et al., 2023). Next, Section 4: Technical Evaluation answers RQ1 by documenting our data-generation and evaluation tasks with 259 CSRs to comparatively analyze Care-Pilot and human-coworker support for a variety of client incivility situations (Fig. 1(a)). Finally, Section 5: User Evaluation answers RQ2 by using Care-Pilot as a technology probe. We conducted a mixed-methods simulation exercise to understand how real CSRs could include it in their client interactions by juxtaposing Care-Pilot usage with CSR’s socio-organizational norms (Fig. 1(b)). Consequently, we contribute:
@@
  • Care-Pilot: an interactive technological artifact to expose CSRs to client incivility and learn healthy, long-term emotional labor practices to improve their own health and support their coworkers.
@@
  • Empirical evidence that LLM–based empathetic support can be engineered to adapt to— and express empathy in— various client-incivility scenarios (RQI). Our results demonstrate that Care-Pilot’s messages are linguistically distinct from both zero-shot approaches and human–coworkers, and moreover, Care-Pilot’s messages were perceived to be more empathetic on several dimensions, including sincerity, actionability, and relatability.
@@
  • End-user insight on the function of LLM–based empathetic support to scaffold them through emotional labor during uncivil interactions (RQ2). Our findings showcase Care-Pilot’s process of redirecting negativity and enhancing CSRs’ self-efficacy. While CSRs envision Care-Pilot addresses important opportunities in workplace social support, they also surfaced the challenges of Care-Pilot emulating holistic human support.
@@
This paper has implications for reimagining how AI-assistants for workers should be designed, and also re-imagining the social norms and policies to accommodate these advancements.
@@
Reflexive Considerations. Front-office work has many stakeholders, including the employer and the clients. However, it is the CSR who bears the burden of repeated emotional labor (Grandey, 2000). The relationship between these stakeholders is asymmetric, as employers can replace personnel, and clients can switch services, but the CSR does not possess the same mobility (Fisk and Neville, 2011). Following from recent works that take a worker-centered perspective (Das Swain et al., 2023a, 2024; Kawakami et al., 2023), our research aims to illuminate challenges faced by disadvantaged workers. This paper focuses on the needs of CSRs and centers their perspective throughout the evaluation. Two authors have experience in front-office roles and direct end-user servicing. Four authors are researchers affiliated with an organization that employs its own CSRs. They helped us access real CSRs (and their resources) to provide feedback on our study design. We recruited participants for this study outside their organization to capture perspectives from different organizational sectors. All evaluations described in the paper were approved by the IRB of the first author’s institute.
@@
(a) Technical evaluation: Compared Care-Pilot’s support messages with those produced by human-coworkers in CSR roles.
@@
(b) User evaluation: Studied participant experiences with Care-Pilot’s emotional support while interacting with uncivil clients.
@@ 2.1. Front-Office Work and Incivility
People who work front-office are the primary human touch-points for direct interaction with clients  [^1] The literature uses the terms “client” and “customer” interchangeably. We use the term “client” to refer to anyone who purchases or avails a good or service.. The end-user, or client, can directly observe and experience the activities of front-office workers (Zomerdijk and de Vries, 2007). Front-office work spans a variety of organizational sectors, including health (e.g., front-desk), hospitality (e.g., air staff), retail (e.g., store associates), and technology (e.g., customer service). For front-office roles, or “pink-collar” work, organizations seek to hire individuals who can emotionally invest in their clients, meet their needs and please them (Hochschild, 2019). Our research focuses on online front-office work, where the employee frequently interacts with their client via text-based communication. This paper refers to these workers as Client Support Representatives or CSRs.
Today, CSRs cater to a variety of clients, and it is common to encounter high-maintenance clients who have unreasonable service demands (Fisk and Neville, 2011). Unfortunately, these often manifest as rude and discourteous behaviors. Koopmann et al. defines client incivility as “low-quality interpersonal treatment that employees receive during service interactions.” Unlike overt verbal abuse, such as name-calling and expletives, or outright violence, incivility is more implicit. According to Andersson and Pearson, the intent to harm is ambiguous and can be deflected, e.g., “I didn’t mean to be rude; I was just in a hurry” (Andersson and Pearson, 1999). Organizational psychologists have noted a possible reason for incivility to be the increase in client entitlement while seeking services (Yagil, 2008) and a lack of consideration for others (Pearson et al., 2000). Meanwhile, a CSR role is often low-wage, considered low-skill, and lacks the decision-latitude needed to respond to mistreatment from clients (Fisk and Neville, 2011). Coupled together, incivility creates a dynamic between the client and CSR, where the latter feels injustice and negative emotion (Fisk and Neville, 2011). Therefore, a CSR’s job involves emotional labor because of two processes; managing the interpersonal demands of interaction and controlling their own negative emotions (Grandey, 2000). Facing client incivility regularly puts a CSR at the risk of experiencing emotional exhaustion (lack of motivation and focus), depersonalization (detachment from others’ emotions), and reduced accomplishment (feeling of low effectiveness) (Maslach et al., 1997). The decline in emotional well-being among CSRs also impacts organizations, leading to increased turnover and a rise in negative work attitudes (Brotheridge and Grandey, 2002). Given the large volume of workers in this role, we need to urgently design solutions for their wellbeing.
@@ 2.2. Social Support and Emotion Regulation
At work, social support is recognized as an important moderator of incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012; Xiao et al., 2022). One kind of social support is emotional support, where a coworker provides sympathy and understanding (Caplan, 1975). This form of support can help reduce work-related stress (House, 1983). Studies have shown that CSRs with more supportive coworkers are able to recover from the negative effects of client incivility, by being able to recognize emotions and guide constructive actions in response (Xu et al., 2014; Wong and Law, 2017). Yet, in practice, CSRs are known to receive low levels of emotional support (Walsh and Deery, 2006) as coworkers can favor client interests over that of an employee (Yang and Lau, 2019). A coworker with high emotional intelligence is someone capable of expressing compassion in a way that helps the CSR cope with negative emotions and proceed to their goal (Law et al., 2004). This behavior of coworkers was the primary metaphor we used to conceive Care-Pilot.
@@
Emotional support can be viewed as a form of emotional coping, or Emotional Regulation (ER). Unlike traditional discussions of ER, coworker support is other-directed (Yang and Lau, 2019). Conversely, CSRs are trained to and expected to regulate their emotions using Surface Acting — they adjust the emotion they present to clients (Grandey, 2000). However, even though Surface Acting is a requirement for roles with emotional labor, suppressing emotions at work can lead to a reduction in job satisfaction (Côté and Morgan, 2002). Instead, Grandey recommends CSRs should engage in Deep Acting — they adjust their perception of the situation. Deep Acting has many strategies, but, CSRs have limited job mobility and do not have the decision latitude to change their clients (Grandey, 2000). As a result, we designed Care-Pilot’s core component Emo-Reframe to achieve ER by helping the CSR cognitively reappraise the negative interaction.
@@
HCI research has explored ER along several dimensions. However, Slovak et al. found that most studies have focused on suppression of emotions (Slovak et al., 2023), and in the CSR context, these methods will have the same limitations as Surface Acting (Côté and Morgan, 2002). Hence, we pursue the ER strategy of cognitive reappraisal or cognitive change. Although, cognitive reappraisal for ER has been explored in HCI in the form of therapy (such as CBT) (Sharma et al., 2023), we find scant evidence for delivering ER in-context and guiding users through the process (Slovak et al., 2023). These gaps motivated us to think beyond training modules and post-hoc support, to implement Care-Pilot for on-task ER within the CSR’s workflow.
2.3. HCI and AI for Emotional Labor
@@
A worker performs emotional labor when their job role expects them to either maintain certain emotions or evoke certain emotions (Hochschild, 1983). While the HCI community has now recognized data work for training responsible AI as a form of emotional labor (Zhang et al., 2024), the predominant attention on this phenomenon has been in the scope of crowd work (Menking and Erickson, 2015) and particularly content moderation (Wohn, 2019; Dosono and Semaan, 2019). Online communities rely on volunteer moderation, and users who participate as moderators often work in a way analogous to front-office work (Section 2.1). Even in an online — primarily text-based — interaction, content moderators have described adopting emotion management techniques to tackle their tasks (e.g., receiving threatening messages from users) (Dosono and Semaan, 2019). Content moderators differ from CSRs, as they are less likely to have synchronous interactions. Having said that, these workers are also subject to reduced emotional wellbeing, such as lack of appreciation and negativity (Wohn, 2019). Another key distinction between content moderation and CSRs is the context of employment. The former is often a voluntary role, whereas the latter is likely to be one’s primary employment. These differences significantly change the socio-organizational dynamics between the worker–client and among the workers. The unique normative structure of emotional labor for CSRs motivated us to conduct a user evaluation to inspect Care-Pilot with actual users.
@@
Since the emotional labor in content moderation is largely composed of harmful content, which can contribute to secondary trauma (Schöpke-Gonzalez et al., 2024), solutions have focused on dynamic content filtering and modulation (Das et al., 2020; Karunakaran and Ramakrishan, 2019). These approaches of preventing or reducing exposure (Steiger et al., 2021) are incompatible with CSR roles where the emotional labor is not caused by explicitly harmful content but by incivility, which is inherently ambiguous, and originates from the client (Andersson and Pearson, 1999).  Cook et al.’s approach to inject positive stimuli in between tasks could be compatible, but their effectiveness was mixed (Cook et al., 2022). Alternatively, Osawa suggested an on-task solution for face-to-face emotion labor, where enhanced glasses make the user’s eyes appear to be emotionally invested (Osawa, 2014). The equivalent of this type of substitution of emotional labor would be the introduction of AI-powered chatbots. Arguably, these chatbots have reinforced client- interactions (Sweezey, 2019), but on one hand, chatbots relieve the task burden of routine inquiries; on the other hand, a large proportion of clients still prefer human CSRs for more complex complaints and better emotional support (Elliott, 2018). Thus, conversational agents do not sufficiently mitigate the challenges of emotional labor on CSR.
@@
3. Care-Pilot: System Description
In traditional front-office work, coworkers are an important social resource. When a CSR encounters uncivil clients, they often consult their coworkers to assuage the excessive emotional labor (Xu et al., 2014; Fisher, 2014). Such consultations can be challenging to interleave into the ongoing task of addressing an aggressive client. Thus, we used OpenAI’s GPT-4o (gpt, 2024) to design an emotional support utility that can be embedded into the CSR’s task environment. Our utility differs from typical conceptions of AI-assistants for work, or “copilots,” that provide informational support by enabling better problem-solving (Cutrona and Russell, 1987), such as coding assistance (Mozannar et al., 2024). Instead, our utility leverages LLMs to provide emotional support by enabling ER through cognitive change (Grandey, 2000). Throughout the paper, we refer to our overarching system for empathetic support as Care-Pilot. This section explains how Care-Pilot was implemented. First, we describe how we harnessed domain knowledge and real complaint datasets to generate uncivil client interactions. Then, we describe how Care-Pilot produces empathetic messages for CSRs who are involved in uncivil interactions. Care-Pilot’s code base primarily relies on LangChain (Topsakal and Akinci, 2023), an open source library and framework to systematically build LLM applications. Particularly, LangChain simplifies the engineering of Chain-of-Thought (CoT) prompts, which are “intermediate natural language reasoning steps” that improve LLM’s ability to tackle commonsense problems (Wei et al., 2022). For instance, our tasks required the LLM to retain memory of the client-CSR conversation, which we achieved via a simple prompt to contextualize history (Fig. A2), and linked it at the beginning of other prompting sequences. Also, throughout our system design, we leveraged descriptions and examples from existing literature. This approach is a form of few-shot learning, where we seed the LLM with prior knowledge to generalize over new tasks (Wang et al., 2020). The data we sample for prompting can be found in the supplementary data and our code is public  [^2] Code repository: https://github.com/vedantswain/care-pilot.git. The following sections elaborate how we prepare this prior knowledge and the appendix provides complete prompt descriptions for adapting our system. Later, in Section 4.4, we compare how Care-Pilot performs against other state-of-the-art LLMs (GPT, Llama, Mistral), which additionally justifies our approach for designing and training Care-Pilot for the purposes of our study.